CPA-CIA in progress provided the following account balances on December 31, 2023 which had been adjusted except for income tax expense:


Cash

....................... ....................... ....................... .....

800,000

Accounts receivable

....................... ....................... ....................... .....

2,800,000

Inventory

....................... ....................... ....................... .....

3,000,000

Property, plant and equipment at carrying amount

....................... ....................... ....................... .....

10,500,000

Accounts payable

....................... ....................... ....................... .....

1,800,000

Income tax payable

....................... ....................... ....................... .....

2,400,000

Bonds payable – due December 31, 2024

....................... ....................... ....................... .....

1,300,000

Share capital

....................... ....................... ....................... .....

2,500,000

Share premium

....................... ....................... ....................... .....

3,000,000

Retained earnings, January 1

....................... ....................... ....................... .....

1,000,000

Net sales and other revenue

....................... ....................... ....................... .....

20,000,000

Costs and expenses

....................... ....................... ....................... .....

12,000,000

Income tax expense

....................... ....................... ....................... .....

2,900,000


The accounts receivable included P1,000,000 due from a customer and payable in quarterly installments of P125,000 under special terms. The last payment is due on December 31, 2025. During the year, estimated tax payment of P500,000 was charged to income tax expense. The income tax rate 30%. Included in property, plant and equipment is land held for sale measured at P1,500,000.


1. On December 31, 2023, what amount should be reported as total current assets?

2. On December 31, 2023, what amount should be reported as total current liabilities?

3. On December 31, 2023, what amount should be reported as total SHE?
Answer #1

1. Computation of the amount that will be shown as current assets on December 31,2023:

Current Assets = Cash + Current Accounts Receivables + Inventory + Land Held for sale

Current Accounts Receivables = P2,800,000 - P1,000,000 + P125,000*4 = P2,300,000

Hence, Current Assets = P800,000 + P2,300,000 + P3,000,000 + P1,600,000 = P7,700,000

2. Computation of the amount that will be shown as current liabilities on December 31,2023:

Current Liabilities = Accounts Payables + Income Tax Payable + Current Bonds Payables

Hence, Current Liabilities = P1,800,000 + P2,400,000 + P1,300,000 = P5,500,000

3. Computation of Shareholder's Equity:

Profit during the year = Net Sale and Revenue - Cost and Expenses - Tax Expenses

Profit during the year = P20,000,000 - P12,000,000 - (P20,000,000 - P12,000,000)*30% = P5,600,000

Shareholder's Equity = Share Capital + Share Premium + Opening Retained Earnings + Profit during the year

Hence, Shareholder's Equity = P1,300,000 + P2,500,000 + P3,000,000 +  P5,600,000 = P12,400,000